# Spectraformer: A Unified Random Feature Framework for Transformer

### v3

@@ 4.3 Analysis of Combinations
In analyzing efficiency, we follow the definition of efficient Transformers in Tay et al. (2022) in terms of memory and computation, the latter interpreted as computational cost. The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF (see Table 6). Hence, when limited computational resource is a concern, PosRF is a better option. Training time and peak memory consumption relies more heavily on component functions than weight matrices, with the exception of FastFoodL. FastFoodL takes significantly less memory consumption at the expense of training time. The accuracy trend for weight matrices is as described above. Table 6 also shows that whilst FastFoodL generally does not work as well due to its low required attention span, its high performance in the task of Text which is low in required attention span and the anomalously performant OPRF-FastFoodL leads to a data skew in the final ranking in Table 3.
Comparing between combinations, we find that for ListOps, OPRF-ORF is the most accurate (38.34%), followed by OPRF-MM (38.08%) and OPRF-QMC (37.69%). OPRF-ORF, being the previous SOTA, is undoubtedly the best choice here since other models maintain the same training time and peak memory consumption with lower accuracy. For Text, the most accurate model is OPRF-SORF (64.76%), followed by SADERF-SORF (64.70%) and PosRF-FastFoodL (64.65%). These all outperform the previous SOTAs (OPRF-ORF and SADERF-ORF) by $4\%+$ whilst maintaining a diversity of training time and peak memory consumption options. Both OPRF-SORF and SADERF-SORF maintain the same level of training time and peak memory consumption as the SOTAs, PosRF-FastFoodL offers much less peak memory consumption at the expense of longer training time. For Retrieval, the most accurate model is SADERF-MM (81.13%), followed by OPRF-MM (81.09%) and SADERF-ORF (81.05%). Whilst these models only perform slightly better than the SOTA SADERF-ORF, OPRF-MM offers less peak memory consumption compared to SADERF-ORF.
5 Conclusion
We present Spectraformer, a framework for approximating and learning the attention kernel in the Transformer. Our paper generalizes past works and presents empirical findings on different component function and weight matrix combinations. We experiment with 18 combinations, of which 14 are novel. Novel combinations perform competitively against, if not more performant than the previous SOTA in the family for LRA textual tasks. We show empirically that the choice of performant weight matrix depends on the required attention span of the task. Our results and the extensibility of our work shows the viability of Spectraformer as a unifying random feature framework of Transformer.
Appendix A Appendix
A.1 Future work, limitations, and broader impact
Spectraformer provides an experimental framework for random features in the Transformer. However, we have not provided formal mathematical proofs for the bounds the these combinations due to the number of combinations discussed. We hope Spectraformer inspires future work to generalizes the proof in such a framework.
Spectraformer has covered a significant number of combinations, however, we have not covered all weight matrices and component functions in the literature. Spectraformer can serve as a foundation for further random feature Transformer research using the techiques not yet experimented with but covered in Appendices A.3, A.4, A.5.
Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning. Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE (Linzen et al., 2018). Transformer-based architecture can also benefit from Spectraformer. Past random feature Transformer works (Likhosherstov et al., 2022; 2023) have experimented with conformer and vision Transformer with ViT. Jeevan & Sethi (2022) provides a good foundation to extend our work. Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification. It could also benefit from ablation studies like the comparison variance in Chowdhury et al. (2022) or in Yao et al. (2023)
Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).s
A.2 Related Work
@@
Spectraformer is based on several theoretical assumptions and choice of mathematical framework and approach. This section discusses alternatives to the choices of Spectraformer. Specifically, Appendix A.2.1 discusses alternative kernelized formulations to Section 2.1; Appendix A.2.2 discusses alternative kernel approximation techniques to random features in Section A.2.2. There are also some component functions and weight matrices which are not covered in the experiments. We introduce them in the related work section in Appendices A.3, A.4, A.5.
A.2.1 Alternative Kernelized Attentions
@@
$\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$ (14)
$\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$ (15)

### v5

@@ 4.5. Efficiency Analysis
In analyzing efficiency, we follow the definition of efficient Transformers in (Tay et al., 2022) in terms of memory and computation, the latter interpreted as computational cost. The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF. Mean accuracy is respectively 53.71, 55.97 and 55.46. Mean time is respectively 1.32, 1.53 and 1.54. Mean memory is 2.31, 2.70, 2.86. Please refer to Table 8. These figures are for 128 random features but with other number of features, the results follow the same trend. SADERF and OPRF having a higher computational requirement than PosRF is due to them having additional computation of the L2 norm of the weight matrix which PosRF does not calculate. Hence, when limited computational resource is a concern, PosRF is a better option. Figure 6 shows that PosRF’s training tmie and memory consumption grows at a significantly lower rate than other methods. Otherwise, OPRF provides the best accuracy. We observe the increase in training time and peak memory consumption as expected due to increasing the number of random features sampled for weight matrices (see Figure 6). Training time and peak memory consumption relies more heavily on component functions than weight matrices, with the exception of FastFoodL. FastFoodL takes significantly less memory consumption at the expense of training time. However, whilst most methods increase training time and memory consumption significantly as the number of features increases, FastFoodL remains constant (see Figure 6).
In addition, Table 5 shows that, in terms of weight matrices, training time and peak memory consumption is relatively stable and does not vary with respect to number of classes, input length, and the required attention span, with the exception for FastFoodL. This is largely due to FastFoodL modifying the attention matrix instead of approximating it. However, we do notice that there is a higher negative correlation between the required attention span and peak memory consumption.
We also analyze efficiency in terms of the number of features. Figure 7 shows that memory usage decrease gradually as the sequence length increases and the training time remains relatively stable.
4.6. Inductive Bias of Efficient Attention
In addition, we also notice that, the top performing models (in Table 1) such as Spectraformer variants, Nystromformer (Xiong et al., 2021), Big Bird (Zaheer et al., 2020), Reformer (Kitaev et al., 2020), Skyformer (Chen et al., 2021) are efficient Transformers that outperform the OT. They can be grouped into the following two induction biases:
  • Low-rank bias (Spectraformer, Nystromformer (Xiong et al., 2021), Skyformer (Chen et al., 2021)) assumes that the true attention matrix is simple, or low rank, in nature.
  • Sparsity bias (Big Bird (Zaheer et al., 2020), Reformer (Kitaev et al., 2020)) assumes that only a small number of attention values in the attention matrix contains meaningful information, ignoring the computation of a significant number of token-to-token attention values, since they only contribute to noise.
Efficient transformers do not just reduce the computational cost, but they also introduce strong induction bias about the structure of the attention matrix. Efficient transformers’ approximation of the OT’s true attention matrix should imply that their performance be upperbounded by the OT’s performance. However, as we have seen in Table 1, they consistently outperform the OT, implying that at least for the characteristics of the LRA, a low-rank bias or a sparsity bias, acting as regularizers on the attention matrix, are helpful induction biases. This aligns with the analysis done in Section 4.4. 128 features will definitely produce a lower rank matrix than 256, 512 features hence models with an even lower-rank bias. This however questions why the OT outperforms 512 and 256-feature models. 128, 256 and 512 feature models have lower rank than the OT and they necessarily produce noise due to the attention approximation process. Having more features likely enlarge the noise in the final approximated attention matrix, making these models fit to spurious noise. Having too few features (64), however, underfits and lead to inexact kernel approximation.
4.7. Method Recommendation
Based on the above experiments, our recommendation is to avoid SORF. SORF performs well on the Text task. However, this is not captured by our metrics (number of output classes, maximum input length, and attention span). There are potential task characteristics that we cannot account for with this exception. Overall, we suggest FastFoodL and QMC among the main weight matrix group of QMC, MM and ORF. They both are highly performant. FastFoodL offers the advantage of low memory consumption at the cost of longer training time, and QMC maintains the average training time and memory consumption.
Our results establish OPRF-FastFoodL as the new state-of-the-art for random feature-based efficient Transformers. More importantly, our findings demonstrate that this class of models is now highly competitive with top-performing methods from other categories like Nystromformer (Xiong et al., 2021) and BigBird (Zaheer et al., 2020). Specifically, OPRF-FastFoodL achieves accuracy on par with BigBird (Zaheer et al., 2020) (57.17% vs. 57.25%) while offering substantial reductions in both peak memory consumption (1.82 GB vs. 4.65 GB) and training time. While Nystromformer attains a higher average accuracy (58.06%), our model provides a valuable alternative with greater memory efficiency (1.82 GB vs. 2.63 GB). These results introduce new, valuable points to the accuracy-memory-time Pareto frontier, offering practitioners compelling options based on their specific performance and resource constraints. In addition to our top-performing model, the Spectraformer framework also produced other novel combinations like OPRF-SGQ and PosRF-MM, which outperform previous RF-based SOTAs (OPRF-ORF and SADERF-ORF) and provide a range of options with different computational profiles. The result mentioned here is available in Table 1.
5. Related Work
@@
Spectraformer is based on several theoretical assumptions various mathematical frameworks. This section discusses alternatives to these assumptions. Specifically, Section 5.1 discusses alternative kernelized formulations to Section 2.1; Section 5.2 discusses alternative kernel approximation techniques to random features in Section 5.2.
5.1. Alternative Kernelized Attentions
@@
(26) $\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$
(27) $\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$